An object is moving along a straight line with a constant velocity of 6 ms-1. The object then speeds up, with constant acceleration, until it reaches a speed of 21 ms-1. During the acceleration phase, the object covers a distance of 134 m. To 2 decimal places, what is the magnitude of the acceleration in ms-2? Give the value only (do not give any units).
Answer by rothauserc(4718)

We use the following equation involving constant acceleration
:
v - vo = a(x-xo)
:
21 - 6 = a(134)
:
a = 15 / 134 = 0.111940299 approx 0.11
: